Let planet A have a mass of 5.44 x 1045 kg and planet B have a mass of 2.77 x 1033 kg. If they are attracted by a gravitational force of 6.59 x 103 N, how far apart are their centers of mass?
Answer #1

Let ,

mass of planet A =m1 = 5.44 × 1045 kg

and Mass of planet B =m2 = 2.77× 1033 kg

Force of atraction = F = 6.59× 1033  N

the distance between their centre of mass will be R.

To find R we use Newton's Law of gravitational attraction,

F= Gm1m2 / R2

where G = Gravitational constant =6.67 × 10-11 m3 kg-1 s-2

there fore ,

R2  = Gm1m2 / F

=  6.67 × 10-11× 5.44 × 1045 × 2.77× 1033 / 6.59× 1033

=100.50× 1067 / 6.59× 1033

R2 =15.25 × 1034

R = 15.25 × 1034

R = 3.91 × 1017 m